OpenStudy (anonymous):

The perimeter of a rectangle is 15 cm, and its dimensions are tripled. What is the perimeter of the larger rectangle?

OpenStudy (anonymous):

20cm 15 cm 45 cm 30 cm

OpenStudy (anonymous):

2*(L+B)=15, now L=3L' and B = 3B' so what is the perimeter?

OpenStudy (anonymous):

what is L and B ????

OpenStudy (anonymous):

Length and breadth respectively.

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

sooo it's 45 cm.

OpenStudy (anonymous):

thanks FFM

OpenStudy (anonymous):

Glad to help sun :)

OpenStudy (anonymous):

P = 2 ( L + w ) = 15 =>P' = 3 * 2 (L + w) = 3 * 15 = 45